Circular No. 6358

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Garnavich and P. Challis, Harvard-Smithsonian Center for Astro-

physics (CfA), on behalf of the High-Z Supernova Search Team, report the discovery of five supernovae on CCD images taken with the Victor Blanco 4-m telescope at Cerro Tololo by A. Clocchiatti, R. Schommer, N. Suntzeff, A. Athey, and E. Turner on Mar. 15 and 16 UT. Spectra of SNe 1996Q, 1996R, 1996T, and 1996U were obtained by R. Kirshner and R. Jayawardhana (CfA) with the Multiple Mirror Telescope (MMT) on Mar. 17 and 18. A spectrum of SN 1996S was obtained by L. Germany and B. Schmidt (Mount Stromlo and Siding Spring Observatories) and R. Stathakis (Anglo-Australian Observatory) with the Anglo-Australian Telescope on Mar. 24. Results derived from the imaging and spectra are as follows:

SN  1996 UT  α_{2000}  δ_{2000}  R  Offset
1996Q Mar. 15 8^{h}46^{m}09.^{s}18 + 0^{o}06′53′′.9 21.8 0′′.6 W, 0′′.3 N
1996R Mar. 15 11 16 10.36 − 0 11 39.3 20.6 2′′.6 E, 6′′.4 S
1996S Mar. 15 11 25 11.16 −35 26 29.1 22.5 0′′.5 W, 1′′.3 N
1996T Mar. 15 10 05 27.66 − 7 27 23.9 21  0′′.0, 0′′.0
1996U Mar. 15 11 50 33.51 +10 26 45.4 22.2 0′′.5 W, 1′′.3 S

The redshifts and estimated supernova types are as follows: SN 1996Q, z = 0.36, ?; SN 1996R, 0.15, Ia; SN 1996S, 0.11, II; SN 1996T, 0.24, Ia; SN 1996U, 0.43, Ia. Narrow emission lines from the host galaxies appear in spectra of SNe 1996Q, 1996T, and 1996U. The redshift of SN 1996R is based on correlating broad features with spectra of well-observed, lowredshift type-Ia supernovae. The MMT spectra also show that an object of magnitude R = 21.0, located 3′′.8 south of SN 1996Q, is a quasi-stellar object at z = 2.06. Finder charts may be requested by sending electronic

mail to pgarnavich@cfa.harvard.edu.

NOVA CASSIOPEIAE 1995 Visual magnitude estimates (cf. IAUC 6331): Mar. 3.13 UT, 9.1 (A.

Hale, Cloudcroft, NM); 9.83, 9.3 (E. Broens, Mol, Belgium); 11.11, 8.9 (W. G. Dillon, Missouri City, TX); 14.80, 9.4 (P. Schmeer, Bischmisheim, Germany); 20.81, 9.4 (S. Baroni, Milan, Italy); 23.96, 8.9 (A. Baransky, Kiev, Ukraine); 24.82, 9.3 (Schmeer).

1996 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green